# Non-monotonic specific entropy on the transition line near the QCD critical point

### v1

@@ Appendix D Topography
We briefly review mathematical tools for exploring the topographic structure of $\hat{s}$ on a general two-dimensional coordinate, $(u,v)=(\mu,\,T)$ or $(r,\,h)$ (see, e.g., [20] for detailed information).
@@
Substituting the fundamental forms (D) into (88), the curvature $\kappa_{n}$ becomes a function of the slope $\lambda={\rm d}v/{\rm d}u$ between two infinitesimally separated points $(u,v)$ and $(u+{\rm d}u,v+{\rm d}v)$:
@@ Appendix E Specific heat
The first and the second terms of (E) correspond to the leading and subleading fluctuations generated in the response to the variance of the $r$ variable, the cross susceptibility $\phi_{r}$ and the specific heat (at constant $h$) $\varepsilon_{r}$, given by (47b) and (47d), respectively. Since the $h$ variable is fixed, the channel with the strongest criticality in Eq. (A), $\phi_{h}$ is absent.
@@ E.2 Specific heat at constant volume
with $\Delta T_{1}$ given below Eq. (III.2).
Let us apply Eq. (110) along the coexistence line by setting $h=\pm 0$ with the susceptibilities given by Eq. (A). Specifically, it describes the slope of the curve with a constant $n$ on the $(T,\hat{s})$ plane at the point where it intersects with the coexistence line. It becomes clear from the analysis in Appendix E.1 that the slopes of curves with constant $n$ and at $h=\pm 0$ may differ at the same point, except at the critical point where both become vertical.

### v2

@@ Appendix D Topography
We briefly review mathematical tools for exploring the topographic structure of $\hat{s}$ on a general two-dimensional coordinate, $(u,v)=(\mu,\,T)$ or $(r,\,h)$ (see, e.g., Ref. [20] for detailed information).
@@
Substituting the fundamental forms (89) into (88), the curvature $\kappa_{n}$ becomes a function of the slope $\lambda={\rm d}v/{\rm d}u$ between two infinitesimally separated points $(u,v)$ and $(u+{\rm d}u,v+{\rm d}v)$:
@@ Appendix E Specific heat
The first and the second terms of (E) correspond to the leading and subleading fluctuations generated in the response to the variance of the $r$ variable, the cross susceptibility $\phi_{r}$ and the specific heat (at constant $h$) $\varepsilon_{r}$, given by (47b) and (47d), respectively. Since the $h$ variable is fixed, the channel with the strongest criticality in Eq. (47), $\phi_{h}$ is absent.
@@ E.2 Specific heat at constant volume
with $\Delta T_{1}$ given below Eq. (27).
Let us apply Eq. (110) along the coexistence line by setting $h=\pm 0$ with the susceptibilities given by Eq. (49). Specifically, it describes the slope of the curve with a constant $n$ on the $(T,\hat{s})$ plane at the point where it intersects with the coexistence line. It becomes clear from the analysis in Appendix E.1 that the slopes of curves with constant $n$ and at $h=\pm 0$ may differ at the same point, except at the critical point where both become vertical.
@@ E.3 At constant pressure
Appendix F Specific entropy maximum on the $(n,\,s)$ plane
Figure 13: Two possible scenarios for the location of the maximum of the specific entropy (green dot) with respect to the critical point (black dot): (a) $\hat{s}_{\rm c}>\cot\alpha_{1}$ and (b) $\hat{s}_{\rm c}<\cot\alpha_{1}$ (see text).
We can also understand the maximum of the specific entropy on the coexistence region boundary in the $(n,\,s)$ plane.
An isentrope with a fixed value of $\hat{s}$ is a straight line passing through the origin. In Fig. 13, we illustrate two isentropes, $s=\hat{s}_{\rm c}n$ (dashed line) and $s=\hat{s}_{\rm max}n$ (green line).
The boundary separating the coexistence region (shaded) from the uniform phases is represented by a black curve on this plane, defined by $h(n,\,s)=\pm 0$, $r<0$ near the critical point.151515 The boundary is a parametric plot, $s=s(r)$ and $n=n(r)$, where the right-hand sides are given by Eqs. (4b) and (4c) with Eq. (11) at $h=\pm 0$. The critical point (black dot) divides the curve into two branches, each serving as the boundary for a different uniform phase. The critical isentrope and the boundary also intersect at another point, the “critical double” (gray dot), as introduced in Fig. 1.
The maximum specific entropy on the coexistence boundary is demonstrated by the existence of an isentrope tangential to the boundary. In Fig. 13, this tangent line and its corresponding tangent point are shown as a green line and a green dot, respectively. It is straightforward to compute the maximum specific entropy by identifying the tangent isentrope from the critical EOS given in Sec. II (see also footnote 15). The results agree with what we discussed in Sec. III on the $(\mu,\,T)$ plane.
Figures 13(a) and 13(b) show the two scenarios where the maximum specific entropy can occur. Each illustrates the maximum on a different branch which can also be classified according to the formula (33). On the $(n,\,s)$ plane, $\hat{s}_{\rm c}$ and $\cot\alpha_{1}$ have a geometric interpretation: the former represents the slope of the critical isentrope (dashed line), while the latter represents the slope of the tangent line to the boundary at the critical point (dotted line). One can show
$\left.\left(\frac{\partial s}{\partial n}\right)_{h}\right|_{{\rm c}}=\cot\alpha_{1}\,,$ (123)
in contrast to the vertical or horizontal slopes on a different plane including one intensive variable, e.g., $(T,\,\hat{s})$ (see Fig. 1) or $(n,\,T)$ diagrams, respectively.
Contour classification becomes more intuitive on the $(n,s)$ plane: straight lines above $s=\hat{s}_{\rm max}n$ do not intersect the boundary (class I); lines between $s=\hat{s}_{\rm max}n$ and $s=\hat{s}_{{\rm c}}n$ intersect only one branch of the boundary (class II); lines below $s=\hat{s}_{{\rm c}}n$ intersect both branches of the boundary (class III).